# Live Music Models

### v1

@@ 2.2 Style embeddings via MusicCoCa
In addition to the two embedding towers, MusicCoCa has a text decoder which can generate audio text captions. In our application this decoder is a shallow 3-layer Transformer which only serves a regularizing purpose.
The MusicCoCa optimization objective, based on the CoCa framework Yu et al. (2022), consists of contrastive and generative loss components which we weigh equally. We train the model using the Adafactor optimizer with a learning rate of $1\times 10^{-4}$ and 1,0001{,}000 warmup steps for a total of 16,00016{,}000 steps. The batch size is $1024$.
@@ 2.3.1 Chunk-based autoregression with coarse context
To achieve RTF $\geq 1\times$, we also propose to use a coarse representation of the audio context. Due to the hierarchical structure of RVQ, lower quantization levels capture the most salient acoustic information. While generating target tokens at the full RVQ depth ($d_{c}=16$) is crucial to maintaining high fidelity, a lower resolution may be sufficient to represent the audio history. Therefore, we use a coarser representation consisting of only the first $4$ RVQ tokens for conditioning on the previous chunks.
@@ 2.4 Encoder-Decoder Language Model
To model this distribution, we use an encoder-decoder Transformer Vaswani et al. (2017) LM trained using T5X Raffel et al. (2020); Roberts et al. (2023). We release pre-trained models using the T5  (Raffel et al., 2020) Base and Large configurations.
@@
A key differentiating factor compared to prior work is our imposed constraint of achieving RTF $\geq 1\times$ generation to enable live interactive applications. Past work such as MusicLM (Agostinelli et al., 2023) proposes a hierarchical cascade of language models to model tokens efficiently, while MusicGen (Copet et al., 2023) proposes a delay pattern—neither approach would achieve RTF $\geq 1\times$ for full bandwidth audio tokens. Instead, similar to Yang et al. (2023); Défossez et al. (2024), our decoder comprises two connected Transformer modules. The “temporal” module constructs a temporal context by processing acoustic frames, where RVQ tokens within each frame are embedded and aggregated to yield a single frame-level embedding. The “depth” module then performs autoregressive prediction of the RVQ indices conditioned on the previous temporal context. With this setup, we achieve RTF=$1.8$ on H100 GPU with the T5 Large configuration.
@@ 3.1 Experimental Set-up
For inference, prompts (text or audio) are embedded and tokenized by MusicCoca using the first $6$ RVQ levels to match training. We sample with classifier-free guidance (CFG) Ho and Salimans (2021), using a temperature of $1.3$, a top-K of $40$, and a CFG weight of $5.0$.
@@ 3.2.2 Generating musical transitions
Figure 2: Prompt transition evaluation. Over $60$ seconds, we transition from embeddings of text prompt A to B by stepwise linear interpolation. Left: Cosine similarity compared to the initial text embedding (blue) and final text embedding (red). Right: Cosine similarity to the interpolation between text embeddings provided to the model. In both plots, lines represent the mean and shaded regions indicate the standard deviation.
@@ 4.1 Style Conditioning via Text and Audio
During inference, we can create a target conditioning vector $\mathbf{c}$ by computing a weighted average of the MusicCoCa embeddings corresponding to $N$ control prompts, provided as either text or audio $\mathbf{c}=\sum_{i=1}^{N}w_{i}M(\mathbf{c_{i}})/\sum_{i}w_{i}$, where $w_{i}$ controls the weight of each prompt. One advantage of using MusicCoCa embeddings, as opposed to attending to text captions, is the ability to perform embedding arithmetic to blend styles—e.g., a weighted sum of embed((techno)) and embed((flute)) gives a good approximation of embed((techno flute))—while also controlling the relative influence of each concept.
Beyond this, using a shared audio-text embedding space for conditioning allows for the use of audio prompts as a more direct way of achieving a specific musical style or instrumentation that may be difficult to express via text. Since audio prompts more closely match the training setting, where style conditioning is obtained from the target audio itself, this type of prompting is also expected to be more effective. Furthermore, we are able to mix multiple audio prompts to achieve interpolations of different styles, or mix combinations of text and audio prompts. Overall, this type of conditioning offers control over high-level characteristics such as genre, style, instrumentation, mood.
@@ D.1 Descriptor-based Conditioning
As seen in Section 4.1, contrastive audio-text embeddings obtained from models such as MusicCoCa and MuLan effectively enable control over the overarching style of generated musical signals. They are not designed, however, to control fine-grained musical attributes. This section investigates the incorporation of additional conditioning features extracted from audio signals using Music Information Retrieval (MIR) methods. A full list of the advanced controls in Lyria RT is provided in Table 3.
@@
We aim at providing a precise control over the tempo of the generated stream, defined using beats-per-minute (BPM). Since we do not have access to BPM annotated audio examples, we use an off-the-shelf beat prediction model Zhao et al. (2022) to annotate our dataset, yielding an estimate for beat positions and an averaged BPM value for the entire track. We start by conditioning our model on the target BPM value rounded to the nearest integer using cross-attention.
@@
In addition to style steering using MuLan embeddings, we aim at providing finer grained controls on the generated stream within a specified musical genre. We use a source separation model Kim et al. (2021) to extract the vocals, bass, drums and other stems from our dataset, and use those stems to generate a set of acoustic descriptors used to further condition our model. We specifically extract peak loudness, spectral centroid and bandwidth, chromas and transients separately for every stem.
@@ D.2 Self-conditioning
Figure 4: Diagram of Lyria RT training / predicting conditioning tokens. Coarse acoustic tokens and quantized MuLan tokens are concatenated and fed to the encoder part of our language model. Control tokens, including BPM, stem balance, brightness, density and chromas (see Section D.1) are predicted first by the decoder, followed by coarse and medium level acoustic tokens. Finely, a small refinement model predicts the fine-scale acoustic tokens as described in Section D.4.
@@
    User defined conditioning might be out of distribution, especially when conditioning tokens are dependent on each other (e.g., natural temporal evolution of a conditioning signal)
@@
    Unconditional generation can be achieved through training-time conditioning dropout, however the resulting samples are empirically worse than those from a model trained unconditionally from scratch
@@ D.3 Constraining Style Embeddings
As mentioned in Section 2.2, we train the language models conditioned on style embeddings computed from the raw audio waveform. This is in contrast to the inference setup where we use text based embeddings. While similar, due to the contrastive training of MuLan, the audio and text embedding spaces are not completely overlapping. Indeed, early experiments showed that a two layer MLP is sufficient to classify embeddings as text-based or audio-based with $90\%$ accuracy. In practice, we notice a non negligible drop in audio quality when predicting from text based embeddings compared to audio embeddings, which is likely due to the mismatch between both embedding spaces.
@@ Appendix E Audio Injection
To allow the user to continuously steer generation via a live audio stream, we propose an audio steering mechanism which we call audio injection. At each generation step, we mix the user’s input audio with the model’s output, tokenize the resulting mix, and feed this as the model’s context for the next generation. This is illustrated in Figure 7. Note, user audio is never played back directly. Rather, the model predicts the continuation of a past context that includes the user audio. Depending on the specific audio injected and the target style, the model may “choose” to repeat or transform the user audio, or may be influenced by its features (dynamics, melody, harmony).

### v2

@@ 2.2 Style embeddings via MusicCoCa
In addition to the two embedding towers, MusicCoCa has a text decoder which can generate audio text captions. In our application this decoder is a shallow 3-layer Transformer which only serves a regularizing purpose. The MusicCoCa optimization objective, based on the CoCa framework Yu et al. (2022), consists of contrastive and generative loss components which we weigh equally. We train the model using the Adafactor optimizer with a learning rate of $1\times 10^{-4}$ and 1,0001{,}000 warmup steps for a total of 16,00016{,}000 steps. The batch size is $1024$.
@@ 2.3.1 Chunk-based autoregression with coarse context
To achieve RTF $\geq 1\times$, we also propose to use a coarse representation of the audio context. Due to the hierarchical structure of RVQ, lower quantization levels capture the most salient acoustic information. While generating target tokens at the full RVQ depth ($d_{c}=16$) is crucial to maintaining high fidelity, a lower resolution may be sufficient to represent the audio history. Therefore, we use a coarser representation consisting of the first $4$ RVQ tokens for conditioning on the previous chunks.
@@ 2.4 Encoder-Decoder Language Model
To model this distribution, we use an encoder-decoder Transformer Vaswani et al. (2017) LM trained with T5X Raffel et al. (2020); Roberts et al. (2023). We release pre-trained models using the T5  (Raffel et al., 2020) Base and Large configurations.
@@
A key differentiating factor compared to prior work is our imposed constraint of achieving RTF $\geq 1\times$ generation to enable live interactive applications. Past work such as MusicLM (Agostinelli et al., 2023) proposes a hierarchical cascade of language models to model tokens efficiently, while MusicGen (Copet et al., 2023) proposes a delay pattern—neither approach would achieve RTF $\geq 1\times$ for full bandwidth audio tokens. Instead, based on McWilliams et al. (2025), our decoder comprises two connected Transformer modules. The “temporal” module constructs a temporal context by processing acoustic frames, where RVQ tokens within each frame are embedded and aggregated to yield a single frame-level embedding. The “depth” module then performs autoregressive prediction of the RVQ indices conditioned on the previous temporal context. With this setup, we achieve RTF=$1.8$ on H100 GPU with the T5 Large configuration.
@@ 3.1 Experimental Set-up
For inference, prompts (text or audio) are embedded and tokenized by MusicCoca using the first $6$ RVQ levels to match training. We sample with classifier-free guidance (CFG) Ho and Salimans (2021); Kreuk et al. (2023); Sanchez et al. (2023), using a temperature of $1.3$, a top-K of $40$, and a CFG weight of $5.0$.
@@ 3.2.2 Generating musical transitions
Figure 2: Prompt transition evaluation. Over $60$s, we transition from embeddings of text prompt A to B by stepwise linear interpolation. Left: Cosine similarity compared to the initial (blue) and final (red) text embedding. Right: Cosine similarity to the interpolation between text embeddings provided to the model. In both plots, lines indicate the mean and shaded regions the standard deviation.
@@ 4.1 Style Conditioning via Text and Audio
During inference, we can create a target conditioning vector $\mathbf{c}$ by computing a weighted average of the MusicCoCa embeddings corresponding to $N$ control prompts, provided as either text or audio $\mathbf{c}=\sum_{i=1}^{N}w_{i}M(\mathbf{c_{i}})/\sum_{i}w_{i}$, where $w_{i}$ controls the weight of each prompt. One advantage of using MusicCoCa embeddings instead of attending to text captions is the ability to perform embedding arithmetic to blend styles—e.g., a weighted sum of embed((“techno”)) and embed((“flute”)) gives a good approximation of embed((“techno flute”))—while also controlling the relative influence of each concept.
Beyond this, a shared audio-text embedding space for conditioning allows for the use of audio prompts as a more direct way of achieving a specific musical style or instrumentation that may be difficult to express via text. Since audio prompts more closely match the training setting, where style conditioning is obtained from the target audio itself, this type of prompting is also expected to be more effective. Furthermore, we are able to mix multiple audio prompts to achieve interpolations of different styles, or mix combinations of text and audio prompts. Overall, this type of conditioning offers control over high-level characteristics such as genre, style, instrumentation, mood.
@@ D.1 Descriptor-based Conditioning
As seen in Section 4.1, contrastive audio-text embeddings obtained from models such as MusicCoCa and MuLan effectively enable control over the overall style of generated musical signals. They are not designed, however, to control fine-grained musical attributes. This section investigates the incorporation of additional conditioning features extracted from audio signals using Music Information Retrieval (MIR) methods. A full list of advanced controls for Lyria RT is provided in Table 3.
@@
We aim to provide precise control over the tempo of the generated stream, defined using beats-per-minute (BPM). Since we do not have access to BPM annotated audio examples, we use an off-the-shelf beat prediction model Zhao et al. (2022) to annotate our dataset, yielding an estimate for beat positions and an averaged BPM value for the entire track. We start by conditioning our model on the target BPM value rounded to the nearest integer using cross-attention.
@@
In addition to style steering using MuLan embeddings, we provide finer grained controls for these features. We use a source separation model Kim et al. (2021) to extract the vocals, bass, drums and other stems from our dataset, and use those stems to generate a set of acoustic features to further condition our model. We specifically extract peak loudness, spectral centroid and bandwidth, chromas and transients separately for every stem.
@@ D.2 Self-conditioning
Figure 4: Diagram of Lyria RT training / predicting conditioning tokens. Coarse acoustic tokens and quantized MuLan tokens are concatenated and fed to the encoder part of our language model. Control tokens, including BPM, stem balance, brightness, density and chromas (see Section D.1) are predicted first by the decoder, followed by coarse and medium level acoustic tokens. Finally, a small refinement model predicts the fine-scale acoustic tokens as described in Section D.4.
@@
    User defined conditioning might be out of distribution, especially when conditioning tokens are dependent on each other (e.g., natural temporal evolution of a conditioning signal).
@@
    Unconditional generation can be achieved through training-time conditioning dropout, however the resulting samples are empirically worse than those from a model trained unconditionally from scratch.
@@ D.3 Constraining Style Embeddings
As mentioned in Section 2.2, we train the language models conditioned on style embeddings computed from the raw audio waveform. This is in contrast to the inference setup where we use text based embeddings. While similar, due to the contrastive training of MuLan, the audio and text embedding spaces are not completely overlapping. Indeed, early experiments showed that a two layer MLP is sufficient to classify embeddings as text-based or audio-based with $90\%$ accuracy. In practice, we notice a non-negligible drop in audio quality when predicting from text based embeddings compared to audio embeddings, which is likely due to the mismatch between both embedding spaces.
@@ Appendix E Audio Injection
To allow the user to continuously steer generation via a live audio stream, we propose an audio steering mechanism that we call audio injection. At each generation step, we mix the user’s input audio with the model’s output, tokenize the resulting mix, and feed this as the model’s context for the next generation. This is illustrated in Figure 7. Note, user audio is never played back directly. Rather, the model predicts the continuation of a past context that includes the user audio. Depending on the specific audio injected and the target style, the model may “choose” to repeat or transform the user audio, or may be influenced by its features (dynamics, melody, harmony).